OpenStudy (anonymous):

Consider the equation xz^2+6yz−2ln(z)=3 as defining z implicitly as a function of x and y. The values of ∂z/∂x and ∂z/∂y at (−3,1,1) are?

OpenStudy (anonymous):

@wio @e.mccormick @Zarkon Any of you capable of helping me on this?

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

\(\dfrac{\partial yz}{\partial x} = y\dfrac{\partial z}{\partial x} + z\dfrac{\partial y}{\partial x}\) using the product rule. This can be significantly simplified.